# Why High-rank Neural Networks Generalize?: An Algebraic Framework with RKHSs

### v1

@@ 5.3 General neural network
If $W$ is not injective, the Koopman operator $K_{W}$ is unbounded. Thus, instead of the standard Koopman operators, we consider weighted Koopman operators. Let $d_{l}\in\mathbb{N}$ and $\Theta_{l}=\{W\in\mathbb{C}^{d_{l}\times d_{l-1}}\}$. For $l=0,\ldots,L-1$, let $\tilde{d}_{l}=\operatorname{dim}(\operatorname{ker}(W_{l+1}))$, $q_{1},\ldots,q_{\tilde{d}_{l}}$ be an orthonormal basis of $\operatorname{ker}(W_{l+1})$, $q_{\tilde{d}_{l}+1},\ldots,q_{d_{l}}$ be an orthonormal basis of $\operatorname{ker}(W_{l+1})^{\perp}$, $\mathcal{X}_{l}=\{\sum_{i=1}^{d_{l}}c_{i}q_{i}\,\mid\,c_{i}\in[a_{i},b_{i}]\}$ for some $a_{i}<b_{i}$ such that $\sigma_{l}(W_{l}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0})\subseteq\mathcal{X}_{l}$, $\mathcal{Y}_{l}=\{\sum_{i=1}^{\tilde{d}_{l}}c_{i}q_{i}\,\mid\,c_{i}\in[a_{i},b_{i}]\}$, and $\mathcal{Z}_{l}=\{\sum_{i=\tilde{d}_{l}+1}^{d_{l}}c_{i}q_{i}\,\mid\,c_{i}\in[a_{i},b_{i}]\}$. Let $W_{l}\sigma_{l-1}(W_{l-1}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\tilde{\mathcal{X}_{l}}\subseteq\mathbb{R}^{d_{l}}$ satisfying $\mu_{\mathbb{R}^{d_{l}}}(\tilde{\mathcal{X}}_{l})>0$. Let $\tilde{\mathcal{H}}_{l}$ and $\mathcal{H}_{l}$ be the same space as in Subsection 5.2, and Let $\eta_{l}(W)=\tilde{K}_{\psi_{l},W}$ be the weighted Koopman operator from $\tilde{\mathcal{H}}_{l}$ to $\mathcal{H}_{l-1}$ with respect to $W$ and $\psi_{l}$, where $\psi_{l}$ is defined as $\psi_{l}(x)=\psi_{l}(x_{1})=1$ for $x\in\mathcal{X}_{l-1}$, where $x=x_{1}+x_{2}$ with $x_{1}\in\mathcal{Y}_{l-1}$ and $x_{2}\in\mathcal{Z}_{l-1}$, and $\psi_{l}(x)=0$ for $x\notin\mathcal{X}_{l-1}$. In addition, let $A_{l}$ be the same operator as in Subsection 5.2. Then, we have
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\det{W_{l}}|_{\operatorname{ker}(W_{l})^{\perp}}|^{-1/2}\leq D}\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})\prod_{l=1}^{L}\mu_{\operatorname{ker}(W_{l})}(\mathcal{Y}_{l-1})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}|_{\operatorname{ker}(W_{l})^{\perp}}|^{1/2}}.$
@@ 5.4 Convolutional neural network
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\beta(\theta_{l})|^{-1/2}\leq D}\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\mu_{\operatorname{ker}(P_{l})}(\hat{\mathcal{Y}}_{l})}{\sqrt{S}\prod_{l=1}^{L}|\beta_{l}(\theta_{l})|^{1/2}}.$
@@ 7 Conclusion and Limitation
In this paper, we derived a new Koopman-based Rademacher complexity bound. Analogous to the existing Koopman-based bounds, our bound describes that neural networks with high-rank weight matrices can generalize well. Existing Koopman-based bounds rely on the smoothness of the function space and the unboundedness of the data space, which makes the result valid for limited neural network models with smooth and unbounded activation functions. We resolved this issue by introducing an algebraic representation of neural network models and constructing an RKHS associated with a kernel defined with this representation. Our bound is valid for a wide range of models, such as those with the hyperbolic tangent, sigmoid, and Leaky ReLU activation functions. Our framework fills the gap between the Koopman-based analysis of generalization bounds and practical situations.
@@ Acknowledgment
This work was supported by JSPS KAKENHI 24K21316, 25H01453, JST BOOST JPMJBY24E2, and JST CREST JPMJCR2015.
@@ Appendix A Proofs
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|}{\sqrt{S}}.$
@@
$=\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|}{\sqrt{S}},$
@@
$\hat{R}(\mathcal{NN}_{c},x_{1},\ldots,x_{S})\leq\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|}{\sqrt{S}}\sup_{|\det{W_{l}}|^{-1/2}\leq D}\prod_{l=1}^{L}|\det W_{l}|^{-1/2}.$
@@
$\qquad\leq\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|}{\sqrt{S}}\sup_{|\det{W_{l}}|^{-1/2}\leq D}|\det W_{1}|^{-1/2}\cdots|\det{W_{L}}|^{-1/2}.$
@@
Theorem 14    Let $\mathcal{F}_{c}=\{F_{c}(\theta_{1},\ldots,\theta_{L},\cdot)\,\mid\,|\det{W_{1}^{*}W_{1}}|^{-1/4},\ldots,|\det{W_{L}^{*}W_{L}}|^{-1/4}\leq D\}$. Assume $p_{c,x}\in\mathcal{K}$ for $x\in\mathcal{X}_{0}$. Let $f_{l}=v\circ W_{L}\circ\sigma_{L-1}\circ\cdots\circ W_{l+1}\circ\sigma_{l}$. Let
for $h\in\tilde{\mathcal{H}}_{l}$. Then, we have
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\sup_{|\det{W_{l}^{*}W_{l}}|^{-1/4}\leq D}\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}^{*}W_{l}|^{1/4}},$
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\det{W_{l}}|_{\operatorname{ker}(W_{l})^{\perp}}|^{-1/2}\leq D}\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})\prod_{l=1}^{L}\mu_{\operatorname{ker}(W_{l})}(\mathcal{Y}_{l-1})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}|_{\operatorname{ker}(W_{l})^{\perp}}|^{1/2}}.$
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\beta(\theta_{l})|^{-1/2}\leq D}\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\mu_{\operatorname{ker}(P_{l})}(\hat{\mathcal{Y}}_{l})}{\sqrt{S}\prod_{l=1}^{L}|\beta_{l}(\theta_{l})|^{1/2}}.$

### v2

@@ 5.3 General neural network
If $W$ is not injective, the Koopman operator $K_{W}$ is unbounded. Thus, instead of the standard Koopman operators, we consider weighted Koopman operators. Let $d_{l}\in\mathbb{N}$ and $\Theta_{l}=\{W\in\mathbb{C}^{d_{l}\times d_{l-1}}\}$. For $l=0,\ldots,L-1$, let $\tilde{d}_{l}=\operatorname{dim}(\operatorname{ker}(W_{l+1}))$, $q_{1},\ldots,q_{\tilde{d}_{l}}$ be an orthonormal basis of $\operatorname{ker}(W_{l+1})$, $q_{\tilde{d}_{l}+1},\ldots,q_{d_{l}}$ be an orthonormal basis of $\operatorname{ker}(W_{l+1})^{\perp}$, $\mathcal{X}_{l}=\{\sum_{i=1}^{d_{l}}c_{i}q_{i}\,\mid\,c_{i}\in[a_{i},b_{i}]\}$ for some $a_{i}<b_{i}$ such that $\sigma_{l}(W_{l}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\mathcal{X}_{l}$, $\mathcal{Y}_{l}=\{\sum_{i=1}^{\tilde{d}_{l}}c_{i}q_{i}\,\mid\,c_{i}\in[a_{i},b_{i}]\}$, and $\mathcal{Z}_{l}=\{\sum_{i=\tilde{d}_{l}+1}^{d_{l}}c_{i}q_{i}\,\mid\,c_{i}\in[a_{i},b_{i}]\}$. Let $W_{l}\sigma_{l-1}(W_{l-1}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\tilde{\mathcal{X}_{l}}\subseteq\mathbb{R}^{d_{l}}$ satisfying $\mu_{\mathbb{R}^{d_{l}}}(\tilde{\mathcal{X}}_{l})>0$.
Figure 3: Construction of $\mathcal{X}_{l}$, $\mathcal{Y}_{l}$, $\mathcal{Z}_{l}$, and $\tilde{X}_{l}$
In this case, to decompose the integral on $\operatorname{ker}(W_{l+1})$ and that on $\operatorname{ker}(W_{l+1})^{\perp}$, we set the orthonormal basis along $\operatorname{ker}(W_{l+1})$ and define $\mathcal{Y}_{l}$ and $\mathcal{Z}_{l}$. Figure 3 schematically shows the construction of $\mathcal{X}_{l}$, $\mathcal{Y}_{l}$, $\mathcal{Z}_{l}$, and $\tilde{X}_{l}$. Let $\tilde{\mathcal{H}}_{l}$ and $\mathcal{H}_{l}$ be the same space as in Subsection 5.2, and Let $\eta_{l}(W)=\tilde{K}_{\psi_{l},W}$ be the weighted Koopman operator from $\tilde{\mathcal{H}}_{l}$ to $\mathcal{H}_{l-1}$ with respect to $W$ and $\psi_{l}$, where $\psi_{l}$ is defined as $\psi_{l}(x)=\psi_{l}(x_{1})=1$ for $x\in\mathcal{X}_{l-1}$, where $x=x_{1}+x_{2}$ with $x_{1}\in\mathcal{Y}_{l-1}$ and $x_{2}\in\mathcal{Z}_{l-1}$, and $\psi_{l}(x)=0$ for $x\notin\mathcal{X}_{l-1}$. In addition, let $A_{l}$ be the same operator as in Subsection 5.2. Then, we have
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\det{W_{l}}|_{\operatorname{ker}(W_{l})^{\perp}}|^{-1/2}\leq D}\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})\prod_{l=1}^{L}\mu_{\operatorname{ker}(W_{l})}(\mathcal{Y}_{l-1})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}|_{\operatorname{ker}(W_{l})^{\perp}}|^{1/2}}.$
@@ 5.4 Convolutional neural network
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\beta(\theta_{l})|^{-1/2}\leq D}\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\mu_{\operatorname{ker}(P_{l})}(\hat{\mathcal{Y}}_{l})}{\sqrt{S}\prod_{l=1}^{L}|\beta_{l}(\theta_{l})|^{1/2}}.$
@@ 7 Conclusion and Limitation
In this paper, we derived a new Koopman-based Rademacher complexity bound. Analogous to the existing Koopman-based bounds, our bound describes that neural networks with high-rank weight matrices can generalize well. Existing Koopman-based bounds rely on the smoothness of the function space and the unboundedness of the data space, which makes the result valid for limited neural network models with smooth and unbounded activation functions. We resolved this issue by introducing an algebraic representation of neural network models and constructing an RKHS associated with a kernel defined with this representation. Our bound is valid for a wide range of models, such as those with the hyperbolic tangent, sigmoid, and Leaky ReLU activation functions. Our framework is the first step to filling the gap between the Koopman-based analysis of generalization bounds and practical situations.
@@ Acknowledgment
This work was supported by JSPS KAKENHI 24K21316, 25H01453, JP24K21316, JST BOOST JPMJBY24E2, and JST CREST JPMJCR2015, JPMJCR25I5, JPMJCR24Q6, JPMJCR24Q1.
@@ Appendix A Proofs
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|{\color[rgb]{0,0,0}E(c)}}{\sqrt{S}}.$
@@
${\color[rgb]{0,0,0}\leq}\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|{\color[rgb]{0,0,0}E(c)}}{\sqrt{S}},$
@@
$\hat{R}(\mathcal{NN}_{c},x_{1},\ldots,x_{S})\leq\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|}{\sqrt{S}}\sup_{|\det{W_{l}}|^{-1/2}\leq D}\prod_{l=1}^{L}|\det W_{l}|^{-1/2}.$
@@
$\qquad\leq\frac{{\color[rgb]{0,0,0}E(c)}\|A_{1}\|\cdots\|A_{L-1}\|\|v\|}{\sqrt{S}}\sup_{|\det{W_{l}}|^{-1/2}\leq D}|\det W_{1}|^{-1/2}\cdots|\det{W_{L}}|^{-1/2}.$
@@
Theorem 15    Let $\mathcal{F}_{c}=\{F_{c}(\theta_{1},\ldots,\theta_{L},\cdot)\,\mid\,|\det{W_{1}^{*}W_{1}}|^{-1/4},\ldots,|\det{W_{L}^{*}W_{L}}|^{-1/4}\leq D\}$. Assume $p_{c,x}\in\mathcal{K}$ for $x\in\mathcal{X}_{0}$. Let $f_{l}=v\circ W_{L}\circ\sigma_{L-1}\circ\cdots\circ W_{l+1}\circ\sigma_{l}$. Let $\alpha(h)=({\int_{W_{l}\mathcal{X}_{l-1}}|h(x)|^{2}d\mu_{\mathcal{R}(W_{l})}(x)}/{\int_{\tilde{\mathcal{X}}_{l}}|h(x)|^{2}\mathrm{d}\mu_{\mathbb{R}^{d_{l}}}(x)})^{1/2}$ for $h\in\tilde{\mathcal{H}}_{l}$. Then, we have
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\sup_{|\det{W_{l}^{*}W_{l}}|^{-1/4}\leq D}\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}^{*}W_{l}|^{1/4}},$
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\det{W_{l}}|_{\operatorname{ker}(W_{l})^{\perp}}|^{-1/2}\leq D}\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})\prod_{l=1}^{L}\mu_{\operatorname{ker}(W_{l})}(\mathcal{Y}_{l-1})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}|_{\operatorname{ker}(W_{l})^{\perp}}|^{1/2}}.$
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})$ $\leq\sup_{|\beta(\theta_{l})|^{-1/2}\leq D}\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\mu_{\operatorname{ker}(P_{l})}(\hat{\mathcal{Y}}_{l})}{\sqrt{S}\prod_{l=1}^{L}|\beta_{l}(\theta_{l})|^{1/2}}.$
@@ B.3 Validity for existing CNN models (LeNet)
Appendix C Comparison of the Koopman-based bounds to existing bounds
Table 1: Comparison of our bound to existing bounds.
Authors Rate Type
Neyshabur et al. (2015) $\frac{2^{L}\prod_{l=1}^{L}\|W_{l}\|_{2,2}}{\sqrt{S}}\rule${0.0pt}{15.0pt} Norm-based
Neyshabur et al. (2018) $\frac{L\max_{l}d_{l}\prod_{l=1}^{L}\|W_{l}\|}{\sqrt{S}}\Big(\sum_{l=1}^{L}\frac{\|W_{l}\|_{2,2}^{2}}{\|W_{l}\|^{2}}\Big)^{1/2}$
Golowich et al. (2018) $\Big(\prod_{l=1}^{L}\|W_{l}\|_{2,2}\Big)\min\bigg\{\frac{1}{S^{1/4}},\sqrt{\frac{L}{S}}\bigg\}$
Bartlett et al. (2017) $\frac{\prod_{l=1}^{L}\|W_{l}\|}{\sqrt{S}}\bigg(\sum_{l=1}^{L}\frac{\|W_{l}^{T}-A_{l}^{T}\|_{2,1}^{2/3}}{\|W_{l}\|^{2/3}}\bigg)^{3/2}$
Wei & Ma (2020) $\frac{(\sum_{l=1}^{L}\kappa_{l}^{2/3}\min\{L^{1/2}\|W_{l}-A_{l}\|_{2,2},\|W_{l}-B_{l}\|_{1,1}\}^{2/3})^{3/2}}{\sqrt{S}}\rule${0.0pt}{16.0pt}
Ju et al. (2022) $\frac{\sum_{l=1}^{L}\theta_{l}\|W_{l}-A_{l}\|_{2,2}}{\sqrt{S}}\rule${0.0pt}{14.0pt}
Li et al. (2021) $\|\mathbf{x}\||\prod_{l=1}^{L}\|W_{l}\|-1|+\gamma_{\mathbf{x}}+\sqrt{\frac{c_{\mathcal{X}}}{S}}\rule${0.0pt}{12.0pt}
Arora et al. (2018) $\hat{r}+\frac{L\max_{i}\|f(x_{i})\|}{\hat{r}\sqrt{S}}\Big(\sum_{l=1}^{L}\frac{1}{\mu_{l}^{2}\mu_{l\rightarrow}^{2}}\Big)^{1/2}\rule${0.0pt}{17.0pt} Compression
Suzuki et al. (2020) $\frac{\hat{r}}{\sqrt{S}}+\sqrt{\frac{L}{S}}\big(\sum_{l=1}^{L}\tilde{r}_{l}(\tilde{d}_{l-1}+\tilde{d}_{l})\big)^{1/2}\rule${0.0pt}{15.0pt}
Hashimoto et al. (2024) $\frac{\|v\|_{H_{L}}}{\sqrt{S}}\prod_{l=1}^{L}\frac{G_{l}\|K_{\sigma_{l}}\|_{H_{l}}\|W_{l}\|^{s_{l-1}}}{\operatorname{det}(W_{l}^{*}W_{l})^{1/4}}\rule${0.0pt}{14.0pt} Koopman-based
Ours $\frac{\|v\|_{\mathcal{L}_{L}}}{\sqrt{S}}\prod_{l=1}^{L}\frac{G_{l}\|K_{\sigma_{l}}\|_{\mathcal{L}_{l}}}{\operatorname{det}(W_{l}^{*}W_{l})^{1/4}}\rule${0.0pt}{14.0pt}
We show the summary of the existing bounds and the proposed bound in Table 1. Here, $\kappa_{l}$ and $\theta_{l}$ are determined by the Jacobian and Hessian of the network $f$ with respect to the $j$th layer and $W_{l}$, respectively. In addition, $\tilde{r}_{l}$ and $\tilde{d}_{l}$ are the rank and dimension of the $j$th weight matrices for the compressed network and $\|\cdot\|_{p,q}$ is the matrix $(p,q)$-norm. We note that although the form of the existing Koopman-based bound and the proposed bound is similar, our bound is applicable to a wider range of deep models, and the factors $G_{l}$ and $\|K_{\sigma_{l}}\|$ are more easily evaluated.
Appendix D Notation table
We provide a notation table 2 that summarizes important notation in the main text.
Table 2: Notation table
$G$ Locally compact group for parameters
$\Theta_{l}$ Set of parameters for the $l$th layer
$L$ Number of layers
$d_{l}$ Width of the $l$th layer
$\mathcal{H}$ Hilbert space for models
$\rho$ Unitary representation of $G$ on $\mathcal{H}$
$K_{\sigma}$ Koopman operator with respect to a function $\sigma$
$W_{l}$ Weight matrix for the $l$th layer
$\sigma_{l}$ Activation function for the $l$th layer
$A_{l}$ Linear operator corresponding to the activation function for the $l$th layer
$f$ Original deep model
$F_{c}$ Regularized model with a parameter $c$
$\mathcal{F}_{c}$ Function class for models
$k$ Positive definite kernel defined as $k((g_{1},\ldots,g_{L}),(\tilde{g}_{1},\ldots,\tilde{g}_{L}))=\left\langle f(g_{1},\ldots,g_{L}),f(\tilde{g}_{1},\ldots,\tilde{g}_{L})\right\rangle_{\mathcal{H}}$
$\phi$ Feature map defined as $\phi(\mathbf{g})=k(\cdot,\mathbf{g})$
$\tilde{\phi}$ Feature map representing models defined as $\tilde{\phi}(\mathbf{g})=f(g_{1},\ldots,g_{L})$, where $\mathbf{g}=(g_{1},\ldots,g_{L})$
$\mathcal{K}$ Hilbert space defined as the closure of $\{\sum_{i=1}^{n}c_{i}\tilde{\phi}(\mathbf{g}_{i})\,\mid\,n\in\mathbb{N},\mathbf{g}_{i}\in G^{L},c_{i}\in\mathbb{C}\}$
$\iota$ Isomorphism that maps $\tilde{\phi}(\mathbf{g})$ to $\phi(\mathbf{g})$